# Nearsightedness in Materials with Indirect Band Gap

### v1

@@ 4. Numerical Experiments
In this section, we present numerical experiments to support our analytical results. We construct a 1D chain with controllable band structure, which allows us to adjust $\mathsf{gap}_{+}$ and $\mathsf{gap}_{-}$.
Let us consider an infinite 1D chain $\Lambda:=\mathbb{Z}$. We construct the tight binding Hamiltonian ${\mathcal{H}}(u)$ by assigning each atom with two atomic orbitals. Only nearest-neighbour interactions are included. For the on-site block of the Hamiltonian matrix, we take a constant diagonal matrix with parameters $c_{1}$ and $c_{2}$; while the off-diagonal block is given by a linear function with a Gaussian envelope:
${\mathcal{H}}({u})_{\ell\ell}=\begin{pmatrix}c_{1}&0\\ 0&c_{2}\end{pmatrix}\quad\text{and}\quad\mathcal{H}({u})_{\ell k}=\begin{pmatrix}f_{1}(r_{\ell k}(u))&f_{3}(r_{\ell k}(u))\\ f_{3}(r_{\ell k}(u))&f_{2}(r_{\ell k}(u))\end{pmatrix}\quad k\neq\ell,$ (4.1)
where the hopping functions $f_{i}(r)=(b_{i}r+a_{i})e^{-d_{i}r^{2}}$ with parameters $a_{i},b_{i},d_{i},i=1,2,3$. One can verify that ${\mathcal{H}}(u)$ satisfies the assumption (TB). Applying the Bloch transform for ${\mathcal{H}}({\bf 0})$ gives
${\mathcal{H}}_{\xi}=\left(\begin{array}[]{cc}c_{1}+2f_{1}(1)\cos(\xi)&2f_{3}(1)\cos(\xi)\\ 2f_{3}(1)\cos(\xi)&c_{2}+2f_{2}(1)\cos(\xi)\end{array}\right),\qquad\forall\xi\in\Gamma^{*}=[-\pi,\pi].$ (4.4)
By solving the eigenvalue problem ${\mathcal{H}}_{\xi}u_{n}(\xi)=\varepsilon_{n}(\xi)u_{n}(\xi)$, we will obtain two energy bands $\{\varepsilon_{n}(\xi)\}_{n=1,2}$, as well as an analytic expression for $\mathsf{gap}_{+}$, which makes it possible for us to build a chain with different $\mathsf{gap}_{-}$ but a roughly fixed $\mathsf{gap}_{+}$; cf. the indirect gap in Figure 1. The computational details can be found in Appendix B.
In practice, we simulate supercell models including 100 atoms for the weak locality test and 200 atoms for the strong locality test, respectively.
Figure 4. 1D chain: Nearsightedness in homogeneous system with fixed $\mathsf{gap}_{+}=2.0$ (left) and $\mathsf{gap}_{-}=0.01$ (right); cf. Section 3.1.
@@
Figure 5. 1D chain: the nearsightedness with local perturbations; cf. Section 3.1
@@
Figure 6. 1D chain: the nearsightedness with global perturbations; cf. Remark 3
In Figure 4, we compare the decay of the density matrix by fixing $\mathsf{gap}_{+}=2.0$ and varying $\mathsf{gap}_{-}=1/2,1/8,1/32,1/128,1/512$, where the decay curves coincide almost perfectly. Moreover, we also compare the results by fixing $\mathsf{gap}_{-}=0.01$ and taking $\mathsf{gap}_{+}=2.0,1.0,0.5,0.25$, which shows a linear dependence on $\mathsf{gap}_{+}$ as displayed in Table 1, strongly supporting our analysis in Section 3.1.
In Figure 5 and Figure 6, we conduct tests with admissible perturbations. Specifically, subfigures (a), (b), and (c) correspond to $\mathsf{gap}_{-}=0.1,0.01$ and $0.001$, respectively. The perturbations have infinity norm of $\varepsilon=10^{-6},10^{-4},10^{-2}$, and $\ell^{2}_{\Upsilon}$ norm of $\varepsilon=10^{-5},10^{-3},10^{-1}$, respectively. The turning points of the curves in each figure shift forward gradually as $\mathsf{gap}_{-}$ decreases and $\varepsilon$ increases. Additionally, we observe that the influence of $\mathsf{gap}_{-}$ becomes increasingly dominant as $\varepsilon$ rises, which aligns perfectly with the analysis presented in Section 3.1.
In Figure 7, we plot the derivative of the density matrix $\big{|}\frac{\partial\rho^{\mathrm{ref}}_{\ell\ell}}{\partial r_{m}}\big{|}$ for $\ell=1$ and $\mathsf{gap}_{-}=1.0,0.1,0.01,0.001$, in both log-log and lin-log scales. We observe that the decay rate slows as $\mathsf{gap}_{-}$ decreases. To further substantiate our conclusions, we provide an example from real systems with a small $\mathsf{gap}_{-}$ in the next section. Additionally, we present the slopes for Figure 7 (b) in Table 2, which displays a square-root dependence on $\mathsf{gap}_{-}$, implying that our results may not be sharp in all situations.
Nevertheless, our results also show that there is a significant pre-asymptotic regime that is not described by our analysis, but which captures relevant scales. We are unaware of a technique to explain the decay in this regime.
(a) log-log scale
(b) lin-log scale
Figure 7. 1D chain: the strong locality in homogeneous system is determined by $\mathsf{gap}_{-}$ rather than by $\mathsf{gap}_{+}$; cf. Proposition 3.2
@@
Table 2. Slopes for the decay curves in Figure 7 (b): the third column displays the optimal fitting between $\mathsf{gap}_{-}$ and slope in a power-law form with parameters $c=-1.51$ and $\alpha=0.51$.
4.2. $\bf{Mg_{2}Si}$
The system we investigate here is an antifluorite-type compounds, a well-known semiconductor used in thermoelectric applications, characterized by a small indirect band gap. We conduct tests using the package $\mathsf{DFTK.jl}$ [17], a library of Julia routines for working with plane-wave density-functional theory algorithms. Although the discussion in the paper focuses on the tight binding model, the property of locality is not limited to this model. Similar results can be found in [1, 18, 30].
The energy bands for ${\rm Mg_{2}Si}$ are presented in Figure 2. The extrema of valence and conduction bands occur at the point $\Gamma$ and $\rm X$ respectively, which results in a small indirect band gap around 0.21 $eV$. However, the $\mathsf{gap}_{+}$ is relatively large around 1.8 $eV$.
We use a supercell model consisting of $4\times 4\times 4$ unit cells to verify the nearsightedness and strong locality. To make the results more convincing, we also present the corresponding tests for $\rm C$. Figures (a) to (d) display the decay in homogeneous systems and the perturbed ones with infinity norm of $\varepsilon=0.01$. The results confirm to a surprising degree that our analysis is applicable to systems of this significantly increased complexity.
@@ 4.2. $\bf{Mg_{2}Si}$
Figure 8. $\rm Mg_{2}Si$ and $\rm C$: the nearsightedness property of homogeneous and perturbed systems; cf. Section 3.1 and Section 3.1.
Finally, we examine the strong locality of $\rm{Mg_{2}Si}$ and $\rm C$ by evaluating the energy Hessian in homogeneous systems, which we computed using the finite difference of site forces $[F(\rho^{\mathrm{ref}})]_{i},i\in\Lambda$, defined by
$d^{2}E_{ij}:=\left|\frac{\partial^{2}E(\rho^{\mathrm{ref}})}{\partial\bm{r}_{i}\partial\bm{r}_{j}}\right|\approx\left|\frac{[F\big{(}\rho(u_{\varepsilon})\big{)}]_{i}-[F(\rho^{\mathrm{ref}})]_{i}}{\partial\bm{r}_{j}}\right|,$ (4.5)
@@
Figure 9. $\rm Mg_{2}Si$ and $\rm C$: the strong locality in a homogeneous system depends on $\mathsf{gap}_{-}$ rather than $\mathsf{gap}_{+}$; cf. Proposition 3.2.
@@ 5. Conclusions
Moreover, we also demonstrated that the impact of perturbations decays as the lattice sites move far away from the perturbations, and this decay is controlled by $\mathsf{gap}_{-}$ and the norm of perturbations. Regarding strong locality (e.g., interatomic interactions), our results, combined with those in [27], indicate that strong locality depends on $\mathsf{gap}_{-}$ and thus may lose fast decay in small gapped systems when considering the Hessian of energy or forces, as shown in Figure 9. However, the results in Figure 7 and Table 2 suggest that, under certain circumstances, the decay exhibits the square-root behaviour with respect to $\mathsf{gap}_{-}$, indicating that our results may not be sharp. A detailed analysis of this effect will be pursued in future work.
@@ Acknowledgements
We gratefully acknowledge stimulating discussions with Antoine Levitt and Alex Watson. HC’s work was supported by the National Natural Science Foundation of China (No. NSFC12371431). JF acknowledges the financial support from the China Scholarship Council (Grant No. 202306040154). CO was supported by NSERC Discovery Grant GR019381 and NFRF Exploration Grant GR022937. JT was supported by the Engineering and Physical Sciences Research Council (EPSRC) Grant EP/W522594/1.
@@ 6. Proofs of the Main Results
6.1. Proof of Section 3.1
In preparation of the proof, we begin by proving a result on perturbations of the spectrum of ${\mathcal{H}}_{\boldsymbol{\xi}}$:
Lemma 6.1 (Perturbations of the Spectrum). Given $\bm{\xi}\in\Gamma^{*}$. Let ${\mathcal{H}}_{\boldsymbol{\xi}}$ be the form of (2.8). There exists a positive constant $c_{\mathrm{r}}$ independent of $\bm{\xi}$, such that
@@
where $c_{\mathrm{r}}=ch_{0}\gamma_{0}^{-(d+1)}$ with $c$ depending on $d,\mathfrak{m},M$ and $N_{\mathrm{b}}$.
@@
For fixed $\ell_{0},k_{0}$, we define the above sum in brackets as the function
@@
which is well-defined and differentiable as $0\leq\beta\leq\frac{1}{2}\gamma_{0}$. Additionally, we see that $F(0)=0$ and the derivative satisfies
@@
$\begin{split}\left\|{\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}-{\mathcal{H}}_{\boldsymbol{\xi}}\right\|_{\mathrm{F}}^{2}=&\sum_{\begin{subarray}{c}\ell_{0},k_{0}\in\Lambda_{0}\\ 1\leq a,b\leq N_{\mathrm{b}}\end{subarray}}\left|\big{[}{\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}\big{]}_{\ell_{0}k_{0}}^{ab}-\big{[}{\mathcal{H}}_{\boldsymbol{\xi}}\big{]}_{\ell_{0}k_{0}}^{ab}\right|^{2}\\ \leq&(N_{\mathrm{b}}h_{0})^{2}\sum_{\begin{subarray}{c}\ell_{0},k_{0}\in\Lambda_{0}\end{subarray}}\left(\sum_{\alpha\in\mathbb{Z}^{d}}e^{-\gamma_{0}|\ell_{0}-k_{0}+\mathsf{A}\alpha|}\big{(}e^{|\ell_{0}-k_{0}+\mathsf{A}\alpha||\bm{\zeta}|}-1\big{)}\right)^{2}\\ \leq&(MN_{\mathrm{b}}h_{0})^{2}\sup_{\begin{subarray}{c}\ell_{0},k_{0}\in\Lambda_{0}\end{subarray}}\left(\sum_{\alpha\in\mathbb{Z}^{d}}e^{-\gamma_{0}|\ell_{0}-k_{0}+\mathsf{A}\alpha|}\big{(}e^{|\ell_{0}-k_{0}+\mathsf{A}\alpha||\bm{\zeta}|}-1\big{)}\right)^{2}\\ \leq&\left(C_{\mathfrak{m},d}MN_{\mathrm{b}}h_{0}\gamma_{0}^{-(d+1)}\right)^{2}|\bm{\zeta}|^{2}\end{split}$ (6.1)
@@
$\mathrm{dist}\left(\sigma\big{(}{\mathcal{H}}_{\boldsymbol{\xi}}\big{)},\,\sigma\big{(}{\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}\big{)}\right)\leq\left\|{\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}-{\mathcal{H}}_{\boldsymbol{\xi}}\right\|_{\mathrm{F}}\leq c_{\mathrm{r}}|\bm{\zeta}|,$ (6.2)
@@
Now we will use Lemma 6.1 to prove Section 3.1:
Proof of Section 3.1.
Let us choose a simple, closed and positively oriented contour $\mathscr{C}$ containing the line segment $i[-R,R]$ for some $R>0$ (chosen later to be sufficiently large) and such that $\varepsilon_{\mathrm{F}}+\mathscr{C}$ encircles the valence bands $\{\varepsilon_{n}\}_{n\leq N_{0}}$ of ${\mathcal{H}}^{\mathrm{ref}}$ and satisfies
$\mathrm{dist}\Big{(}\varepsilon_{\mathrm{F}}+\mathscr{C},\sigma({\mathcal{H}}^{\mathrm{ref}})\Big{)}>0.$ (6.3)
Applying the contour integral representation in (3.1), we have
[…]
where $\nu:\Gamma^{*}\rightarrow\mathbb{C}$ is a mapping to be determined, satisfying $\varepsilon_{N_{0}}(\bm{\xi})<\nu(\bm{\xi})<\varepsilon_{N_{0}+1}(\bm{\xi})$ for all $\bm{\xi}\in\Gamma^{*}$. For simplicity, let us define $R_{z}(\bm{\xi}):=\big{(}z-{\mathcal{H}}_{\boldsymbol{\xi}}\big{)}^{-1}$ and note
$\widehat{\rho}(\bm{\xi})\coloneqq\oint_{\mathscr{C}+\nu(\bm{\xi})}(z-{\mathcal{H}}_{\boldsymbol{\xi}})^{-1}\frac{\mathrm{d}z}{2\pi i}=\oint_{\mathscr{C}+\nu(\bm{\xi})}R_{z}(\bm{\xi})\,\frac{\mathrm{d}z}{2\pi i}\quad\in\mathbb{C}^{MN_{\mathrm{b}}\times MN_{\mathrm{b}}},$ (6.5)
@@
$\rho^{\mathrm{ref}}_{\ell k,ab}=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.70335pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.59167pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.10834pt}}\!\int_{\Gamma^{*}}\big{[}\widehat{\rho}(\bm{\xi})\big{]}_{\ell_{0}k_{0},ab}\,e^{i\bm{\xi}\cdot(\ell-k)}\mathrm{d}\bm{\xi},$ (6.6)
@@
We first define $\mathsf{g}(\bm{\xi})\coloneqq\varepsilon_{N_{0}+1}(\bm{\xi})-\varepsilon_{N_{0}}(\bm{\xi})$ and note that $\mathsf{gap}_{+}=\min_{\Gamma^{\star}}\mathsf{g}$. Next, introduce $\alpha>0$ to be determined later, and
$D_{\mathrm{R}}\coloneqq\left\{\bm{\xi}+i\bm{\zeta}\colon\bm{\xi},\bm{\zeta}\in\mathbb{R}^{d},|\bm{\zeta}|\leq\min\Big{\{}\tfrac{1}{2}\gamma_{0},\,\tfrac{\alpha}{c_{\mathrm{r}}}\mathsf{gap}_{+}\Big{\}}\right\}$ (6.7)

### v2

@@ 4. Numerical Experiments
In this section, we present numerical experiments to support our analytical results. Specifically, we demonstrate the dependence of nearsightedness on $\mathsf{gap}_{+}$ and its stability under perturbations, as well as the strong locality dependence on $\mathsf{gap}_{-}$. We also include experiments that validate the potential sharp estimates. To this end, we construct 1D toy models with tunable band structures, allowing us to adjust $\mathsf{gap}_{+}$ and $\mathsf{gap}_{-}$ in a controlled manner. In addition, we also illustrate our results on a realistic 3D system, $\rm Mg_{2}Si$.
@@
Let us consider an infinite 1D chain $\Lambda:=\mathbb{Z}$. We construct the tight binding Hamiltonian ${\mathcal{H}}(u)$ by assigning each atom with two atomic orbitals. Only nearest-neighbour interactions are included.
We suppose that ${\mathcal{H}}(u)$ is given by ${\mathcal{H}}(u)_{\ell k}=h\big(\bm{r}_{\ell k}(u)\big)$ with
(4.1) $h\big(\bm{r}_{\ell k}(u)\big)$ $=\begin{cases}\bm{c}&\text{if }\ell=k\\ \bm{f}\big(r_{\ell k}(u)\big)&\text{if }\ell\not=k\end{cases}$
and $f_{ij}(r)\coloneqq(a_{ij}+b_{ij}r)\eta(r)$ for parameters $\bm{a},\bm{b},\bm{c}\in\mathbb{R}^{2\times 2}_{\rm sym}$ and for some smooth cut off function $\eta$ with $\eta=1$ on $[0,1]$ and $\eta=0$ on $2,\infty)$. One may verify that ${\mathcal{H}}(u)$ satisfies (TB) with Bloch transform of ${\mathcal{H}}(0)$ given by
(4.2) ${\mathcal{H}}_{\xi}=\bm{c}+2\bm{f}\cos(\xi),\quad\text{with }\xi\in\Gamma^{\star}=[-\pi,\pi].$
where $\bm{f}\coloneqq\bm{f}(1)$. By solving the eigenvalue problem ${\mathcal{H}}_{\xi}u_{n}(\xi)=\varepsilon_{n}(\xi)u_{n}(\xi)$, we obtain two energy bands $\varepsilon_{\pm}(\xi)$, as well as an analytic expression for $\mathsf{gap}_{+}$, which enables us to build a chain with varying $\mathsf{gap}_{-}$ while keeping $\mathsf{gap}_{+}$ approximately fixed; cf. the vertical gap in Figure 1.
In the following, we will consider two “types” of band gaps: (i) first, we consider $c_{12}=c_{21}=0$ and take the off-site block to be a full matrix. We also consider (ii) the case when the on-site block is approximately diagonal ($c_{12}\approx 0$), while the off-site block is diagonal (that is, $f_{12}=f_{21}=0$) and thus $\varepsilon_{\pm}(\xi)$ resemble cosine-like bands. The following figures illustrate band structures of the two types Hamiltonian under various $\mathsf{gap}_{-}$ and $\mathsf{gap}_{+}$ settings.
(i) In Figure 4 (a) and (b), we fix $\mathsf{gap}_{+}=1.0$ and let $\mathsf{gap}_{-}$ decease from $1.0$ to $0.1$, showing an indirect gap, with the curvature norm $m_{\infty}$ uniformly bounded. This means we would expect $\sqrt{\mathsf{gap}_{-}}$ dependence in the derivatives of $\rho(u)$, as predicted by Theorem 3 (c.f. Figure 10). Notably, as $\mathsf{gap}_{+}\to 0$, $m_{\infty}$ diverges, see Figure 4 (c).
(a) $\mathsf{gap}_{+}=\mathsf{gap}_{-}=1.0$
(b) $\mathsf{gap}_{+}=1.0,~\mathsf{gap}_{-}=0.1$
(c) $\mathsf{gap}_{+}=0.25,~\mathsf{gap}_{-}=0.01$
Figure 4. Band structure of type (i) model.
(ii) Figure 5 displays the cosine-like bands with a direct band gap. As $\mathsf{gap}_{+}\to 0$, $m_{\infty}$ remains uniformly bounded and one expects a $\sqrt{\mathsf{gap}_{+}}$ dependence in the nearsightedness estimate as predicted by Theorem 3 (cf. Figure 9).
(a) $\mathsf{gap}_{+}=\mathsf{gap}_{-}=1.0$
(b) $\mathsf{gap}_{+}=\mathsf{gap}_{-}=0$
Figure 5. Band structure of type (ii) model.
The specific examples and computational details can be found in Appendix B.
4.2. Nearsightedness
We conduct the tests using the type (i) model under two settings: one with fixed $\mathsf{gap}_{+}$ and varing $\mathsf{gap}_{-}$, and the other with fixed $\mathsf{gap}_{-}$ and varing $\mathsf{gap}_{+}$, to validate the dependence on $\mathsf{gap}_{+}$. In practice, we simulate supercell models including 200 atoms.
In Figure 6, we compare the decay of the density matrix by fixing $\mathsf{gap}_{+}=2.0$ and varying $\mathsf{gap}_{-}$ across the values $1/2,1/8,1/32,1/128,1/512$, where the decay curves align almost perfectly. Moreover, we also compare the results by fixing $\mathsf{gap}_{-}=0.01$ and taking $\mathsf{gap}_{+}=2.0,1.0,0.5,0.25$, which shows a linear dependence on $\mathsf{gap}_{+}$ as displayed in Table 1.
Figure 6. 1D chain: nearsightedness in homogeneous system with (a) fixed $\mathsf{gap}_{+}=2.0$ and (b) $\mathsf{gap}_{-}=0.01$; cf. Section 3.
@@ 4.2. Nearsightedness
In Figure 7 and Figure 8, we conduct tests with admissible perturbations. Specifically, subfigures (a), (b) and (c) correspond to $\mathsf{gap}_{-}=0.1,0.01$ and $0.001$, respectively. The perturbations have infinity norm of $\varepsilon=10^{-6},10^{-4},10^{-2}$, and $\ell^{2}_{\Upsilon}$ norm of $\varepsilon=10^{-5},10^{-3},10^{-1}$, respectively. The turning points of the curves in each figure shift forward gradually as $\mathsf{gap}_{-}$ decreases and $\varepsilon$ increases. Additionally, we observe that the influence of $\mathsf{gap}_{-}$ becomes increasingly dominant as $\varepsilon$ rises, which aligns perfectly with the analysis in Section 3.
@@
Figure 7. 1D chain: nearsightedness with local perturbations; cf. Section 3
@@
Figure 8. 1D chain: nearsightedness with global perturbations; cf. Remark 3
In Figure 9, we conduct the tests using the type (ii) model, which yields cosine-like bands for suitable parameter choices; see Appendix B for details. We plot the decay curves of the density matrix for $\mathsf{gap}_{+}(=\mathsf{gap}_{-})=0.1,0.01,0.001$ on a lin-log scale. The $\mathsf{gap}_{+}$-slope relation in Figure 9(b) exhibits a square-root dependence on $\mathsf{gap}_{+}$.
$\mathsf{gap}_{+}(=\mathsf{gap}_{-})$ slope $c\mathsf{gap}_{+}^{\alpha}$
0.1 -0.0585 -0.0589
0.01 -0.0208 -0.0205
0.001 -0.0071 -0.0071
Figure 9. 1D chain: the sharp estimate for nearsightedness in homogeneous system; the third column in (b) is the optimal fitting between $\mathsf{gap}_{+}$ and slopes with c=0.17c=-0.17, $\alpha=0.47$; cf. Theorem 3.
4.3. Strong locality
In the following experiments, we focus exclusively on the type (i) model, which allows for more controlled tuning of $\mathsf{gap}_{-}$ with fixed $\mathsf{gap}_{+}$, enabling clearer observation of the decay behaviour.
In Figure 10, we plot the derivative of the density matrix $\big|\frac{\partial\rho^{\mathrm{ref}}_{\ell\ell}}{\partial r_{m}}\big|$ for $\ell=1$ and $\mathsf{gap}_{-}=1.0,0.1,0.01,0.001$, in lin-log scales. We observe that the decay becomes slower as $\mathsf{gap}_{-}$ decreases. Additionally, we present the slopes for Figure 10(a), which displays a square-root dependence on $\mathsf{gap}_{-}$.
@@ 4.3. Strong locality
Figure 10. 1D chain: the strong locality in homogeneous system; the third column displays the optimal fitting between $\mathsf{gap}_{-}$ and slope with c=1.51c=-1.51 and $\alpha=0.51$; cf. Theorem 3
To further substantiate our conclusions, we provide an example from real systems with a small $\mathsf{gap}_{-}$ in the next section.
$4.$ $\bf{Mg_{2}Si}$
The system we investigate here is an antifluorite-type compounds, a well-known semiconductor used in thermoelectric applications, characterized by a small indirect band gap. We conduct tests using the package $\mathsf{DFTK.jl}$ [18], a library of Julia routines for working with plane-wave density-functional theory algorithms. Although the discussion in the paper focuses on the tight binding model, the property of locality is not limited to this. Similar results can be found in [1, 19, 33].
The energy bands for $\rm Mg_{2}Si$} are presented in Figure 2. The extrema of valence and conduction bands occur at the point $\Gamma$ and $\rm X$ respectively, which results in a small $\mathsf{gap}_{-}$ around 0.21 $eV$. However, the $\mathsf{gap}_{+}$ is relatively large around 1.8 $eV$.
We use a supercell model consisting of $4\times 4\times 4$ unit cells to verify the nearsightedness and strong locality. To make the results more convincing, we also present the corresponding tests for $\rm C$. In Figure 11, (a) to (d) display the decay in homogeneous systems and the perturbed ones with infinity norm of $\varepsilon=0.01$.
@@
Figure 11. $\rm Mg_{2}Si$ and $\rm C$: the nearsightedness property of homogeneous and perturbed systems; cf. Section 3 and Section 3.
@@
Figure 12. $\rm Mg_{2}Si$ and $\rm C$: the strong locality in a homogeneous system depends on $\mathsf{gap}_{-}$ rather than $\mathsf{gap}_{+}$; cf. (3.8).
Finally, we examine the strong locality of $\rm{Mg_{2}Si}$ and $\rm C$ by evaluating the energy Hessian in homogeneous systems, which we computed using the finite difference of site forces $F(\rho^{\mathrm{ref}})]_{i},i\in\Lambda$, defined by
(4.3) $d^{2}E_{ij}:=\left|\frac{\partial^{2}E(\rho^{\mathrm{ref}})}{\partial\bm{r}_{i}\partial\bm{r}_{j}}\right|\approx\left|\frac{[F\big(\rho(u_{\varepsilon})\big)]_{i}-[F(\rho^{\mathrm{ref}})]_{i}}{\partial\bm{r}_{j}}\right|,$
@@ 5. Conclusions
Moreover, we also demonstrated that the impact of perturbations decays as the lattice sites move far away from the perturbations, and this decay is controlled by $\mathsf{gap}_{-}$ and the norm of perturbations. Regarding strong locality (e.g., interatomic interactions), our results, combined with those in [28], indicate that strong locality depends on $\mathsf{gap}_{-}$ and thus may lose fast decay in small gapped systems when considering the Hessian of energy or forces, as shown in Figure 12. Moreover, the results in Figure 9 and Figure 10 suggest that, under certain circumstances, the decay exhibits the square-root behaviour with respect to $\mathsf{gap}_{-}$ or $\mathsf{gap}_{+}$, indicating the possibility of sharp estimates. To explore this further, we conducted a focused case study to support and justify these observations. A more detailed analysis of this effect will be pursued in future work.
@@ Acknowledgements
We gratefully acknowledge stimulating discussions with Antoine Levitt and Alexander Watson. HC’s work was supported by the National Natural Science Foundation of China (No. NSFC12371431). JF acknowledges the financial support from the China Scholarship Council (Grant No. 202306040154). CO was supported by NSERC Discovery Grant GR019381 and NFRF Exploration Grant GR022937. JT was supported by the Engineering and Physical Sciences Research Council (EPSRC) Grant EP/W522594/1.
@@ 6. Proofs of the Main Results
6.1. Proof of Section 3
In preparation of the proof, we begin by proving a result on perturbations of the spectrum of ${\mathcal{H}}_{\boldsymbol{\xi}}$, given by (2.8):
Lemma 6.1 (Perturbations of the Spectrum). Fix $\bm{\xi}\in\Gamma^{*}$. Then,
@@ 6.1. Proof of Section 3
where $c_{\mathrm{r}}=ch_{0}\gamma_{0}^{-(d+1)}$ with $c>0$ depending only on $d,\mathfrak{m},M$ and $N_{\mathrm{b}}$.
@@
For fixed $\ell_{0},k_{0}$, we define
@@
which is well-defined and differentiable on $\beta\in[0,\frac{1}{2}\gamma_{0}]$. Additionally, we see that $F(0)=0$ and the derivative satisfies
@@
(6.1) $\begin{split}\left\|{\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}-{\mathcal{H}}_{\boldsymbol{\xi}}\right\|_{\mathrm{F}}^{2}=&\sum_{\begin{subarray}{c}\ell_{0},k_{0}\in\Lambda_{0}\\ 1\leq a,b\leq N_{\mathrm{b}}\end{subarray}}\left|\big[{\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}\big]_{\ell_{0}k_{0}}^{ab}-\big[{\mathcal{H}}_{\boldsymbol{\xi}}\big]_{\ell_{0}k_{0}}^{ab}\right|^{2}\\ \leq&(N_{\mathrm{b}}h_{0})^{2}\sum_{\begin{subarray}{c}\ell_{0},k_{0}\in\Lambda_{0}\end{subarray}}\left(\sum_{\alpha\in\mathbb{Z}^{d}}e^{-\gamma_{0}|\ell_{0}-k_{0}+\mathsf{A}\alpha|}\big(e^{|\ell_{0}-k_{0}+\mathsf{A}\alpha||\bm{\zeta}|}-1\big)\right)^{2}\\ \leq&(MN_{\mathrm{b}}h_{0})^{2}\sup_{\begin{subarray}{c}\ell_{0},k_{0}\in\Lambda_{0}\end{subarray}}\left(\sum_{\alpha\in\mathbb{Z}^{d}}e^{-\gamma_{0}|\ell_{0}-k_{0}+\mathsf{A}\alpha|}\big(e^{|\ell_{0}-k_{0}+\mathsf{A}\alpha||\bm{\zeta}|}-1\big)\right)^{2}\\ \leq&\left(C_{\mathfrak{m},d}MN_{\mathrm{b}}h_{0}\gamma_{0}^{-(d+1)}\right)^{2}|\bm{\zeta}|^{2}\end{split}$
@@
(6.2) $\mathrm{dist}\left(\sigma\big({\mathcal{H}}_{\boldsymbol{\xi}}\big),\,\sigma\big({\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}\big)\right)\leq\left\|{\mathcal{H}}_{\bm{\xi}+i\bm{\zeta}}-{\mathcal{H}}_{\boldsymbol{\xi}}\right\|_{\mathrm{F}}\leq c_{\mathrm{r}}|\bm{\zeta}|,$
@@
Now we will use Lemma 6.1 to prove Section 3:
Proof of Section 3.
Let us choose a simple, closed and positively oriented contour $\mathscr{C}$ containing the line segment i[R,R]i[-R,R] for some $R>0$ (chosen later to be sufficiently large) and such that $\varepsilon_{\mathrm{F}}+\mathscr{C}$ encircles the valence bands $\{\varepsilon_{n}\}_{n\leq N_{0}}$ of ${\mathcal{H}}^{\mathrm{ref}}$ and satisfies
(6.3) $\mathrm{dist}\Big(\varepsilon_{\mathrm{F}}+\mathscr{C},\sigma({\mathcal{H}}^{\mathrm{ref}})\Big)>0.$
Applying the contour integral representation in (3), we have
(6.4) $\rho^{\mathrm{ref}}_{\ell k,ab}$ $=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.91281pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.30664pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.13547pt}}\!\int_{\Gamma^{*}}\bigg[\oint_{\mathscr{C}+\nu(\bm{\xi})}(z-{\mathcal{H}}_{\boldsymbol{\xi}})^{-1}\frac{\mathrm{d}z}{2\pi i}\bigg]_{\ell_{0}k_{0},ab}e^{i\bm{\xi}\cdot(\ell-k)}\mathrm{d}\bm{\xi},$
where $\nu:\Gamma^{*}\rightarrow\mathbb{C}$ is a mapping to be determined, satisfying $\varepsilon_{N_{0}}(\bm{\xi})<\nu(\bm{\xi})<\varepsilon_{N_{0}+1}(\bm{\xi})$ for all $\bm{\xi}\in\Gamma^{*}$. For simplicity, let us define $R_{z}(\bm{\xi}):=\big(z-{\mathcal{H}}_{\boldsymbol{\xi}}\big)^{-1}$ and
(6.5) $\widehat{\rho}(\bm{\xi})\coloneqq\oint_{\mathscr{C}+\nu(\bm{\xi})}(z-{\mathcal{H}}_{\boldsymbol{\xi}})^{-1}\frac{\mathrm{d}z}{2\pi i}=\oint_{\mathscr{C}+\nu(\bm{\xi})}R_{z}(\bm{\xi})\,\frac{\mathrm{d}z}{2\pi i}\quad\in\mathbb{C}^{MN_{\mathrm{b}}\times MN_{\mathrm{b}}}.$
@@
(6.6) $\rho^{\mathrm{ref}}_{\ell k,ab}=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.91281pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.30664pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.13547pt}}\!\int_{\Gamma^{*}}\big[\widehat{\rho}(\bm{\xi})\big]_{\ell_{0}k_{0},ab}\,e^{i\bm{\xi}\cdot(\ell-k)}\mathrm{d}\bm{\xi},$
@@
We first define $\mathsf{g}(\bm{\xi})\coloneqq\varepsilon_{N_{0}+1}(\bm{\xi})-\varepsilon_{N_{0}}(\bm{\xi})$ and note that $\mathsf{gap}_{+}=\min_{\Gamma^{*}}\mathsf{g}$. Next, introduce $\alpha>0$ to be determined later, and
(6.7) $D_{\mathrm{R}}\coloneqq\left\{\bm{\xi}+i\bm{\zeta}\colon\bm{\xi},\bm{\zeta}\in\mathbb{R}^{d},|\bm{\zeta}|\leq\min\Big\{\tfrac{1}{2}\gamma_{0},\,\tfrac{\alpha}{c_{\mathrm{r}}}\mathsf{gap}_{+}\Big\}\right\}$